OpenStudy (anonymous):

tan A=1|2 and tan B =1|3 then the value of A+B is

OpenStudy (anonymous):

Use tan sum formula to express tan(A+B) in terms of Tan A and Tan B and then A+B is arctan that.

OpenStudy (anonymous):

\[A+B=\tan^{-1}((\tan A+\tan B )/(1-\tan A*\tan B))\] tanA=1/2,tanB=1/3 \[A+B=\tan^{-1} (5/6*6/5)\] \[A+B=\Pi/4\] i hope u unterstood

OpenStudy (anonymous):

Oh, I see, u just want answers...